# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v3

@@ G.2.2 Astex & PoseBusters RMSD results
In Figures 12 and 13, we report the ligand RMSD values of each baseline method across the Astex Diverse and PoseBusters Benchmark datasets, with relaxation being applied in the context of the PoseBusters Benchmark dataset. In short, we see that most methods are relatively similar in terms of their ligand RMSD distributions, with RoseTTAFold-All-Atom and our ensembling consensus baseline (i.e., Ensemble (Con)), however, offering more condensed distributions overall. Interestingly, for Astex Diverse, TULIP also appears to produce a uniquely confined ligand RMSD distribution.
G.3 Expanded CASP15 results
G.3.1 Overview of expanded results
@@
In this section, we begin by reporting additional CASP15 benchmarking results in terms of each baseline method’s multi-ligand RMSD and lDDT-PLI distributions as violin plots. Subsequently, we report successful ligand docking success rates as well as RMSD and lDDT-PLI results specifically for the single-ligand CASP15 targets. Lastly, we report all the above single and multi-ligand results specifically using only the CASP15 targets for which the ground-truth (experimental) structures are publicly available, to enable reproducible future benchmarking and follow-up works.
G.3.2 Multi-ligand RMSD and lDDT-PLI
@@
To start, Figures 14 and 15 report each method’s multi-ligand RMSD and lDDT-PLI distributions with and without relaxation. We see that NeuralPLexer and Ensemble (Con) produce the most tightly bound and accurate RMSD and lDDT-PLI distributions overall.
G.3.3 All single-ligand results
@@
Next, Figures 16, 17, 18, and 19 display each method’s single-ligand CASP15 docking success rates, PoseBusters validity rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) DiffDock-L and Ensemble (Con) are the only methods capable of successfully docking any single-ligand CASP15 complexes. (2) AutoDock Vina produces the most PB-valid single-ligand complexes overall, with DiffDock-L shortly behind. (3) DiffDock-L, AutoDock Vina, and Ensemble (Con) appear to achieve the most tightly bound and accurate RMSD distributions. (4) In contrast to (3), only Ensemble (Con) appears to achieve top results in terms of lDDT-PLI compared to the other baseline methods.
G.3.4 Single and multi-ligand results for public targets
@@
Lastly, for completeness and reproducibility, Figures 20, 21, 22, and 23 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 24, 25, 26, and 27 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes. However, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes.

### v4

@@ H.2.2 Astex & PoseBusters RMSD results
In Figures 17 and 18, we report the ligand RMSD values of each baseline method across the Astex Diverse and PoseBusters Benchmark datasets, with relaxation being applied in the context of the PoseBusters Benchmark dataset. In short, we see that most methods are relatively similar in terms of their ligand RMSD distributions, with RoseTTAFold-All-Atom and Ensemble (Con), however, offering more condensed distributions overall. Interestingly, for Astex Diverse, TULIP also appears to produce a uniquely confined ligand RMSD distribution.
H.3 Expanded CASP15 results
H.3.1 Overview of expanded results
@@
In this section, we begin by reporting additional CASP15 benchmarking results in terms of each baseline method’s multi-ligand RMSD and lDDT-PLI distributions as violin plots. Subsequently, we report successful ligand docking success rates as well as RMSD and lDDT-PLI results specifically for the single-ligand CASP15 targets. Lastly, we report all the above single and multi-ligand results specifically using only the CASP15 targets for which the ground-truth (experimental) structures are publicly available, to support reproducible future benchmarking and follow-up works.
H.3.2 Multi-ligand RMSD and lDDT-PLI
@@
To start, Figures 19 and 20 report each method’s multi-ligand RMSD and lDDT-PLI distributions with and without relaxation. We see that NeuralPLexer and Ensemble (Con) produce the most tightly bound and accurate RMSD and lDDT-PLI distributions overall.
H.3.3 All single-ligand results
@@
Next, Figures 21, 22, 23, and 24 display each method’s single-ligand CASP15 docking success rates, PoseBusters validity rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) DiffDock-L and NeuralPLexer are the only DL methods capable of successfully docking any single-ligand CASP15 complexes. (2) AutoDock Vina produces the most PB-valid single-ligand complexes overall, with TULIP shortly behind. (3) DiffDock-L and AutoDock Vina appear to achieve the most tightly bound and accurate RMSD distributions. (4) In contrast to (3), only DiffDock-L-Vina appears to achieve top results in terms of lDDT-PLI compared to the other baseline methods.
H.3.4 Single and multi-ligand results for public targets
@@
Lastly, for completeness and reproducibility, Figures 25, 26, 27, and 28 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 29, 30, 31, and 32 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes, since once again NeuralPLexer achieves top results in this multi-ligand context. However, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes. In short, we find that DiffDock-L-Vina performs the best in this setting.